OpenStudy (anonymous):

Determine algebraically whether the function is even, odd, or neither even nor odd. f(x) = -3x4 - 2x - 5 (note: the 3x4 is an exponent as in 3x to the 4th power)

OpenStudy (anonymous):

^4 is the usual ascii notation for an exponent

OpenStudy (anonymous):

but good clarification nonetheless

OpenStudy (anonymous):

ah my bad

OpenStudy (anonymous):

if f(-1) = f(1), its even if f(-1) = - f(1), its odd otherwise its neither

OpenStudy (anonymous):

for example, x^2 is an even function (-1)^2 = (1)^2 .. its even x^3 is an odd function (-1)^3 = - (1)^3

OpenStudy (anonymous):

so.. this would be an odd function?

OpenStudy (anonymous):

let test it out ... f( 1) = -3 - 2 - 5 = -10 f(-1) = -3 + 2 - 5 = -6 is f(-1) = f(1) ? -6 = -10 is f(-1) = - f(1) ? -6 = 10 of neither, then neither

OpenStudy (anonymous):

So its neither? jesus im so confused XD where did the -10 and -6 come from

OpenStudy (anonymous):

when x = -1, f(-1) = -10 its just working the function, replace x by -1 and add up the terms :/

OpenStudy (anonymous):

if we try this algebrically, we would use a more generic input like x = -a, and x=a

OpenStudy (anonymous):

the -1 and 1 are just simpler values to "see" with

OpenStudy (anonymous):

f( a) = -3a^4 - 2a - 5 f(-a) = -3a^4 + 2a - 5 -3a^4 - 2a - 5 ?=? -3a^4 + 2a - 5 -3a^4 - 2a ?=? -3a^4 + 2a - 2a ?=? 2a -a ?=? a; only if a = 0

OpenStudy (anonymous):

okay so the answer to the problem would be neither, correct? i just need an answer at this point im terrible with functions, i get the concept, its just applying it in pre calculus is an issue for me.

OpenStudy (anonymous):

neither, yes

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y+%3D++-3x4+-+2x+-+5 visually, if the graph is a mirror image across the y axis, its even if its a rotation thru the origin by 180 degrees, its odd